OpenStudy (anonymous):

Linear equation (solve by substitution) 16x+19y=82 6x+y=190 This is the step I think of next... -19(6x+y=190)

OpenStudy (nikato):

So you are trying to substitute the second equation into the first?

OpenStudy (nikato):

And I think you made a typo somewhere in your first equation

OpenStudy (anonymous):

I am trying to make the y's match up so I can get rid of them (one has to be negative one has to be positive) This way they cancel than I can move on to find y.

OpenStudy (nikato):

Oh ok. Then yes, that would be your next step. I got confused because this method is called elimination, not substitution

OpenStudy (anonymous):

My paper says substitution. But from how it is I guess it can be both. Ok well after that step I do this 16x+19y=82 + -114x+-19y=-3610

OpenStudy (nikato):

Correct. And now add them up

OpenStudy (anonymous):

-98x=-3610

OpenStudy (nikato):

You forgot to add the right hand side So -98x= -3528 And now solve for x

OpenStudy (anonymous):

yeah i just noticed and its 36

OpenStudy (anonymous):

Oh my god -_____- I just realized I got it wrong every time because i forgot the other side of the =

OpenStudy (nikato):

Haha, it's ok. So what did you get for y after plugging 36 into x for one of the equations

OpenStudy (anonymous):

for y im getting -34.6 but she said there are no decimals or fractions

OpenStudy (nikato):

I don't think you did it right. Try it again. You should get -26

OpenStudy (anonymous):

So I have to make it negative by subtracting 576?

OpenStudy (nikato):

This is pluggin into second equation 6(36)+y=190 216+y=190 -216. -216 ---------------- y=-26

OpenStudy (anonymous):

Im doing the wrong problem -____- I cannot think today

OpenStudy (nikato):

Or if you were doing the first, 16(36)+19y=82 576+19y=82 -576. -576 ------------------- 19y= -494 --- ----- 19. 19 y=-26

OpenStudy (anonymous):

Thank you so much! I feel so retarded.

OpenStudy (nikato):

No problem. Wait till you here this. If you actually follow the directions, substitution would have been much faster. Subtract 6x on both sides in your second equation so you have y= 190-6x Substitute this into the first equation so 16x + 19(190-6x)=82 And now solve for x And then do the same thing as the previous problem to solve for y

OpenStudy (anonymous):

-_____- I swear my teacher doesn't give us the right notes with titles and instructions...

OpenStudy (nikato):

Haha. Sounds like you have a great teacher

OpenStudy (anonymous):

WONDERFUL!!!